# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

### v1

@@ 2. Stratifications of the lower $p$-series
Being $p$-adic analytic, the pro-$p$ group $G$ has finite rank. By [7, Prop. 3.9], there exists $j\in\mathbb{N}$ such that $U=P_{j}(G)$ is uniformly powerful of rank $\mathrm{rk}(U)=\dim(G)=d$ and such that $P_{i}(G)$ is powerfully embedded in $U$ for all $i\in\mathbb{N}$ with $i\geq j$. We recall that there is an explicit isomorphism of categories translating between uniformly powerful pro-$p$ groups and powerful $\mathbb{Z}_{p}$-Lie lattices; see [7, Sec. 4 and 9] and [20]. This means that the underlying set of the pro-$p$ group $U$ can be equipped in a canonical way with the structure of a $\mathbb{Z}_{p}$-Lie lattice $L$ carrying the same topology and incorporating essentially all information about $U$. In particular, the construction is such that exponentiation in $U$ corresponds to scalar multiplication in $L$ and that the conjugation action of $G$ on $U$ translates to a continuous $\mathbb{Z}_{p}$-linear action of $G$ on $L$. In order to distinguish between the two parallel points of view, it is convenient to write $\underline{x}$ for $x\in U$ when it features as an element of the Lie lattice $L$ rather than as a group element. Using this notation, the earlier statements can be phrased as follows: the (identity) map $U\to L$, $x\mapsto\underline{x}$ is a homeomorphism; we have $\underline{x^{a}}=a\underline{x}$ for all $x\in U$, $a\in\mathbb{Z}_{p}$; and conjugation, encapsulated in the canonical homomorphism $G\to\operatorname{Inn}(G)$, induces a continuous homomorphism
@@
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [20, Thm. 1.4].
Clearly, $j^{*}=0$ gives $L_{j^{*}}=L=\lambda_{j^{*}}(L)$, and we contend that, in fact, $L_{i^{*}}=\lambda_{i^{*}}(L)$ for all $i\geq j$. Let $i\in\mathbb{N}$ with $i>j$. Since $P_{i-1}(G)$ is powerful, its Frattini subgroup equals $P_{i-1}(G)^{p}$ and corresponds to the Lie sublattice $pL_{i^{*}-1}$. Multiplicative cosets of $P_{i-1}(G)^{p}$ in $P_{i-1}(G)$ are the same as additive cosets of $pL_{i^{*}-1}$ in $L_{i^{*}-1}$, and multiplication in $P_{i-1}(G)/P_{i-1}(G)^{p}$ is the same operation as addition $L_{i^{*}-1}/pL_{i^{*}-1}$; see [7, Cor. 4.15]. Since
@@
In conjunction with [8, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [7, Sec. 4] and recorded in [20, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
@@
where we write $[\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [10, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [20, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
@@
is a subset of $\check{U}_{i}$. Furthermore, [7, Prop. 3.7 and Sec. 4.2] show that $\check{V}_{i}$ is the disjoint union of $p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}$ cosets of the open normal subgroup
@@
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [20, Prop. A.2] and [6, Lem. 1.9.4]. This gives
@@
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [7, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\leq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
@@ 3. Hausdorff spectra with respect to the lower $p$-series
We write $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [22, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
@@
Then we see that, for all sufficiently large $i\in\mathbb{N}$,
$\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert=\lfloor i\xi_{j(r)}\rfloor-m$
@@
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [7, Prop. 3.9]) with corresponding Lie sublattice $M\leq_{\mathrm{o}}L$, then
@@
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\leq U_{i^{*}}$, and accordingly $p^{m}L\leq L_{i^{*}}$. By [7, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence

### v2

@@ 2. Stratifications of the lower $p$-series
Being $p$-adic analytic, the pro-$p$ group $G$ has finite rank. By [5, Prop. 3.9], there exists $j\in\mathbb{N}$ such that $U=P_{j}(G)$ is uniformly powerful of rank $\mathrm{rk}(U)=\dim(G)=d$ and such that $P_{i}(G)$ is powerfully embedded in $U$ for all $i\in\mathbb{N}$ with $i\geq j$. We recall that there is an explicit isomorphism of categories translating between uniformly powerful pro-$p$ groups and powerful $\mathbb{Z}_{p}$-Lie lattices; see [5, Sec. 4 and 9] and [15]. This means that the underlying set of the pro-$p$ group $U$ can be equipped in a canonical way with the structure of a $\mathbb{Z}_{p}$-Lie lattice $L$ carrying the same topology and incorporating essentially all information about $U$. In particular, the construction is such that exponentiation in $U$ corresponds to scalar multiplication in $L$ and that the conjugation action of $G$ on $U$ translates to a continuous $\mathbb{Z}_{p}$-linear action of $G$ on $L$. In order to distinguish between the two parallel points of view, it is convenient to write $\underline{x}$ for $x\in U$ when it features as an element of the Lie lattice $L$ rather than as a group element. Using this notation, the earlier statements can be phrased as follows: the (identity) map $U\to L$, $x\mapsto\underline{x}$ is a homeomorphism; we have $\underline{x^{a}}=a\underline{x}$ for all $x\in U$, $a\in\mathbb{Z}_{p}$; and conjugation, encapsulated in the canonical homomorphism $G\to\operatorname{Inn}(G)$, induces a continuous homomorphism
@@
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [15, Thm. 1.4].
Clearly, $j^{*}=0$ gives $L_{j^{*}}=L=\lambda_{j^{*}}(L)$, and we contend that, in fact, $L_{i^{*}}=\lambda_{i^{*}}(L)$ for all $i\geq j$. Let $i\in\mathbb{N}$ with $i>j$. Since $P_{i-1}(G)$ is powerful, its Frattini subgroup equals $P_{i-1}(G)^{p}$ and corresponds to the Lie sublattice $pL_{i^{*}-1}$. Multiplicative cosets of $P_{i-1}(G)^{p}$ in $P_{i-1}(G)$ are the same as additive cosets of $pL_{i^{*}-1}$ in $L_{i^{*}-1}$, and multiplication in $P_{i-1}(G)/P_{i-1}(G)^{p}$ is the same operation as addition in $L_{i^{*}-1}/pL_{i^{*}-1}$; see [5, Cor. 4.15]. Since
@@
In conjunction with [6, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [5, Sec. 4] and recorded in [15, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
@@
where we write $[\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [8, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [15, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
@@
is a subset of $\check{U}_{i}$. Furthermore, [5, Prop. 3.7 and Sec. 4.2] show that $\check{V}_{i}$ is the disjoint union of $p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}$ cosets of the open normal subgroup
@@
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [15, Prop. A.2] and [4, Lem. 1.9.4]. This gives
@@
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [5, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\leq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
@@ 3. Hausdorff spectra with respect to the lower $p$-series
We write $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [17, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
@@
For each $i\in\mathbb{N}$ the order of the finite cyclic group $(H_{2}+L_{i})/L_{i}$ equals the order of $y_{r}$ modulo $L_{i}$. Thus we see that, for all sufficiently large $i\in\mathbb{N}$,
$\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert=\max\{\lfloor i\xi_{k}\rfloor-v_{p}(a_{r,k})\mid 1\leq k\leq j(r)\}=\lfloor i\xi_{j(r)}\rfloor-m$
@@
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [5, Prop. 3.9]) with corresponding Lie sublattice $M\leq_{\mathrm{o}}L$, then
@@
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\leq U_{i^{*}}$, and accordingly $p^{m}L\leq L_{i^{*}}$. By [5, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence